Assume that you believe that a certain stock will face a high volatility in the near future, since the company will release their annual report in two weeks time, and you think that it will contain some surprises. You don't know if it will be good or bad surprises,why you create a straddle by purchasing a call and a put option with the same strike price 106. The put option premium is 3.8 and the call option premium is 8. What is the profit of this portfolio if the stock price at the time of expiration is 119?
Answer #1

Straddle Profit=MAX(119-106,0)+MAX(106-119,0)-3.8-8=1.2

Profit per share=1.20
Profit per contract=1.20*100=120